OpenStudy (anonymous):

Solve the following equation....plz show work. (5x - 1)/2 + (-3x - 1)/4 = 9/2

OpenStudy (chaise):

Best off cross multiplying once again, to be honest.

OpenStudy (chaise):

\[\frac{5x-1}{2}+\frac{-3x-1}{4}=\frac{9}{2}\] \[\frac{4(5x-1)+2(-3x-1)}{8}=\frac{9}{2}\] Multiply by 8 both sides and expand the brackets.

OpenStudy (chaise):

20x-4-6x-2=36 I think. 14x-6=36 14x=42 x=42/14 x=3

OpenStudy (anonymous):

5x−1/2+−3x−1/4=9/2 4(5x−1)+2(−3x−1)/8=9/2 20x-4-6x-2=64/2 20x-6x=32+4+2 14x=38 x:38/14 x=2.71

OpenStudy (anonymous):

Which one is correct.

OpenStudy (anonymous):

i think chaise's is right

OpenStudy (anonymous):

I think?

OpenStudy (anonymous):

sugargurl's answer ,there is something wrong 8*9=72 not 64

OpenStudy (anonymous):

how did chaise get 36 then?

OpenStudy (anonymous):

she is wrong

OpenStudy (anonymous):

So, there both wrong?

OpenStudy (anonymous):

sorry !i made a mistake

OpenStudy (anonymous):

8*9/2=36

OpenStudy (anonymous):

chaise's is right

OpenStudy (anonymous):

okay thanks.